# Operations & Supply Chain Management: Principles and Practice

### v1

@@ 3 Supply Chain Management
To a significant extent, the major strands of stakeholder research in OSCM reflect the schism between normative and instrumental stakeholder thinking. To date, the most prominent strand of stakeholder OSCM research focuses on the roles played by stakeholders in supporting or inhibiting sustainable outcomes throughout firms’ operations and supply chains. Stakeholders are understood to exert pressure on firms for sustainable business conduct, thereby suggesting that firms pay attention to improving their environmental and social conditions (Carter and Liane Easton, 2011; Meixell and Luoma, 2015). This attention extends beyond individual firms’ operations, in that it also includes their direct (Foerstl et al., 2010) and indirect (Hartmann and Moeller, 2014) suppliers. When firms fail to meet sustainability expectations, they risk reputational damage and legal issues (Bregman et al., 2015). This creates what is referred to as supply chain sustainability risk (SCSR) – a condition within the supply chain that can provoke adverse stakeholder reactions (Hofmann et al., 2014). In the era of global sourcing and pervasive information, SCSR presents a significant challenge (Busse et al., 2017). Research has indicated that successful sustainable supply chain management (SSCM) requires rethinking stakeholder relationships. Firms should view non-governmental organisations and other non-traditional actors as collaborative partners rather than adversaries (Pagell and Wu, 2009). An attentive and cooperative stance helps firms address SCSR by incorporating diverse stakeholder knowledge and expectations (Wu et al., 2014). Stakeholders differ in their roles, interests, and degrees of influence on the economic, environmental, and social dimensions (Meixell and Luoma, 2015). For instance, employees and customers may wish for a firm’s success, while competitors and the media may be indifferent or even oppositional (Hofmann et al., 2014). Firms must avoid treating stakeholders as homogeneous, and instead dedicate specific attention to distinct groups (Gualandris et al., 2015).
Regarding stakeholder-OSCM research in the instrumental tradition, trust, collaboration, and mutual investment are seen as key to fostering innovation and operational performance. Studies have consistently linked supply chain collaboration to improved outcomes (Co and Barro, 2009). Long-term buyer-supplier relationships enhance coordination, whereas strategic partnerships increase competitiveness (Choi and Hartley, 1996; Carr and Pearson, 1999). Collaborative behaviours such as shared planning and flexibility are strongly associated with trust (Johnston et al., 2004). Trust and commitment promote efficiency, productivity, and effectiveness, requiring influential advocates to foster cooperation (Morgan and Hunt, 1994). Additionally, mutual trust enables firms to adopt innovative practices, reduce transaction costs, and maintain stronger partnerships, thus further cementing their competitive edge. Supply chain visibility is crucial for effective stakeholder management because it enables traceability and transparency across supply chain processes (Tse and Tan, 2012). However, firms often lack visibility beyond their first-tier suppliers, particularly in fast-moving industries such as retail and fashion (Roth et al., 2008). Limited visibility hampers supply chain risk management, leading to knowledge gaps, loss of control, and heightened risk (Durach et al., 2015; Busse et al., 2017). Addressing this challenge involves expanding the scope of supply chain boundaries and integrating diverse actors (Carter et al., 2015). For instance, many firms struggle to gain a full view of their upstream supply chains, which makes it difficult to identify potential risks and inefficiencies. Enhanced transparency not only mitigates risks, but also fosters greater accountability and trust among stakeholders, which is increasingly critical in today’s interconnected world.
Considerable research has begun to examine the structural, attitudinal, and informational precursors of successful stakeholder management. Carter et al. (2015) argued that OSCM research generally neglected important supply chain actors not directly involved with the material flows within supply chains, and that supply chain boundaries are typically unclear. This visibility boundary poses a severe management problem, since “beyond the visible range, the agent has no choice but to accept what happens there” (Carter et al., 2015, p. 90). Prior research has acknowledged that missing visibility in supply chains is a critical factor in effective supply chain risk management (Durach et al., 2015). Lack of visibility may cause knowledge deficits, loss of control, and distrust, thereby enhancing sustainability and corporate social risks (Busse et al., 2017).
@@
Similarly, research suggests that firms can adopt various stances toward their stakeholders, ranging from adversarial to welcoming (Pagell and Wu, 2009). Pagell and Wu (2009) observed that leading firms reconceptualise who is in the supply chain, such that they regard not only their direct buyers and suppliers as part of the supply chain, but also other stakeholders. Essentially, they suggest the opening up of firms to these stakeholders. Further studies have also found that some firms leverage the expertise and skills of stakeholders, resulting in better-informed managerial decision-making (Wong et al., 2015). Wong et al. (2015, p. 56) argued that “feedback from [$\dots$] stakeholders represents key resources because (they) sometimes know more about the environmental problems facing part of the supply chains than the focal firm”. Stakeholders can provide assistance, develop policies, engage in evaluation and monitoring, and identify improvement potential in a firm’s upstream supply chain with regard to sustainability (Gualandris et al., 2015; Wong et al., 2015). In the context of SCSR, leading firms have begun to proactively search for valuable information to help them identify their SCSR (or other objectives) by constantly scanning the external environment or conducting regular stakeholder consultations and round tables (Foerstl et al., 2010).
@@
Recent macro-environmental uncertainty and turbulence has encourage OSCM scholarship to explore the role of stakeholders in supporting organisational success in light of these conditions. Indeed, Freeman (1984) himself drew attention to “the quantity and kinds of change that occur in the business environment” (Freeman, 1984, , chapter 5), necessitating a new paradigm of strategic management. The growth and diversity of stakeholder-related research in OSCM provide scholars with numerous sources for further reading and inspiration. We have drawn the reader’s attention to some important review articles. For example, Aaltonen and Kujala (2016) review literature on stakeholder management in the project management field; Bhattacharya and Fayezi (2021) review research on stakeholder theory in supply chain management; de Gooyert et al. (2017) review stakeholder management research in the operations research field; Govindan and Bouzon (2018) examine stakeholder-related research in reverse logistics. Sustainable supply chain management has been reviewed extensively by Carter and Liane Easton (2011) and Sarkis et al. (2011). Additionally, there are a number of excellent reviews on the development of stakeholder management outside OSCM, including Laplume et al. (2008). Technology, especially AI and big data analytics, and government policy play a significant role in shaping how firm-stakeholder relationships evolve and are managed, and rapid changes in these in recent years will continue to show that stakeholder relationships play an increasingly salient role in OSCM.
@@
The bullwhip effect refers to a supply chain phenomenon where the variance of customer demand is amplified in the production (or supplier) orders. Figure 5 highlights the weekly demand from a grocery retailer for a UK dairy product. The production order variance is 7.68 times that of demand. This amplification is not uncommon in real supply chains. The highly variable demand placed on the upstream echelons of a supply chain creates excess costs. For example, highly variable demand requires extra capacity to meet peak demands; extra inventory is needed to protect service levels; reduced transport efficiency; hiring, firing, and onboarding costs increase.
@@
Control theory has long been one of the most powerful ways of solving linear bullwhip problems. Simon (1952) applied the Laplace to a production control problem described by differential equations. Towill (1982) studied an inventory replenishment policy denoted IOBPCS (inventory and order-based production control system) in continuous time using the Laplace transform. John et al. (1994) added a work-in-progress feedback loop to the IOBPCS to create the APIOBPCS model (automatic pipeline, inventory, and order-based production control system). Continuous-time bullwhip measures are commonly determined by the system’s noise bandwidth.
Most supply chains operate on discrete time, where one ordering decision per planning period is made. Vassian (1955) studied a discrete-time replenishment system, finding the inventory optimal order-up-to (OUT) policy with z-transforms. Disney (2001) used the z-transform to study the dynamics of a vendor-managed inventory supply chain. The Fourier transform can be used to understand how supply chains create bullwhip under arbitrary (stationary and non-stationary) demands and (constant) lead times by investigating the system’s frequency response. Dejonckheere et al. (2003) revealed the OUT policy with exponential smoothing and moving average forecasts always creates bullwhip, for all demand patterns and lead times. Li et al. (2014) extended these results, showing the OUT policy with Holt’s forecasts always creates bullwhip, but the OUT policy with damped trend forecasts does not always create bullwhip for certain demand processes. Linear systems can also be studied directly in the time domain and using difference equations and expectation (Hosoda and Disney, 2012). Larger, more complex linear supply chain models can also be analyzed using state space techniques and matrix methods (Gaalman and Disney, 2006).
@@
Surveys and secondary data can also be used to investigate bullwhip. At an individual product level, bullwhip appears almost inevitable. However, when measured at the company level, perhaps aggregated over a financial reporting period of many months, bullwhip may not exist (Cachon et al., 2007). Bullwhip has also been studied empirically: Fransoo and Wouters (2000) consider how to measure bullwhip in real supply chains (which may have convergent and/or divergent material flows). Disney et al. (2013) report on a project removing bullwhip from a global printer toner supply chain. Meng et al. (2018) improve the bullwhip behaviour in a UK dairy product manufacturer.
@@
Most bullwhip studies consider high-volume demand where the items are ordered and replenished every planning period. Planning periods range from several hours to a day in grocery supply chains, or weekly (most likely) to monthly in manufacturing environments. With no ordering (capacity-related) costs, the OUT policy is the optimal replenishment policy for minimizing inventory holding and backlog costs (Karlin and Scarf, 1958). When the objective function includes bullwhip-related (capacity-related) costs, the OUT policy is no longer optimal, Boute et al. (2022); the dual base stock (DBS) policy is known to be the optimal policy (Boute and Van Mieghem, 2015; Gijsbrechts et al., 2024). However, the DBS policy is not a demand replacement policy. The DBS policy has a region of inaction where demand is not fully replaced and only $k$ units (where $k$ is the nominal capacity without overtime) are ordered. Generally, only numerical solutions for the two base stock levels in the DBS policy can be obtained (Gijsbrechts et al., 2024).
The high-performing linear proportional (POUT) policy, grounded in control theory, is amenable to analytical study. The POUT policy generalises the OUT policy by incorporating two proportional feedback controllers into the policy, an established technique in the design of hardware systems. The feedback controllers in the POUT policy regulate the speed at which discrepancies in inventory and WIP levels are corrected (Disney and Towill, 2003). While the POUT policy requires careful set-up to avoid instability (Disney, 2008), it is the optimal linear quadratic regulator for minimising costs proportional to the variance of order and inventory (Wang and Disney, 2016). When piece-wise linear convex inventory costs and capacity costs associated with an installed capacity and overtime are present Boute et al. (2022) show the POUT performs well compared to the optimal non-linear DBS policy.
@@
Most bullwhip studies assume demand is random. The simplest random demand is the independent and identically distributed (i.i.d.) demand. Forecasted with minimum mean squared error (MMSE) forecasts, i.i.d. demand produces a bullwhip ratio of unity, indicating the OUT policy acts as pass-on-orders policy.
@@
Closed loop supply chains are popular areas for bullwhip studies (Ponte et al., 2022). Marketplace returns can have a positive effect on the production of new items. However, care has to taken as it is not always economical to remanufacture everything returned, even when the remanufacturing cost is less than the cost to produce new items (Hosoda et al., 2021).
Finally, issues with global dual sourcing have been studied from a bullwhip perspective. The tailored-base-surge allocation scheme (Allon and Van Mieghem, 2010), places large constant orders to the offshore supplier, whereas the nearshore supplier satisfies peak demands with a shorter lead time. Boute et al. (2022) showed that local SpeedFactories could maintain tighter inventory control than a single offshore supplier, facilitating the relocation of production back home, before reaching price parity.

### v2

@@ 3 Supply Chain Management
To a significant extent, the major strands of stakeholder research in OSCM reflect the schism between normative and instrumental stakeholder thinking. To date, the most prominent strand of stakeholder OSCM research focuses on the roles played by stakeholders in supporting or inhibiting sustainable outcomes throughout firms’ operations and supply chains. Stakeholders are understood to exert pressure on firms for sustainable business conduct, thereby suggesting that firms pay attention to improving their environmental and social conditions (Carter and Liane Easton, 2011; Meixell and Luoma, 2015). This attention extends beyond individual firms’ operations, in that it also includes their direct (Foerstl et al., 2010) and indirect (Hartmann and Moeller, 2014) suppliers. Stakeholder pressures increase awareness of sustainability in the supply chain, push buying firms to adopt sustainability-related goals, and influence them in implementing sustainability in the supply chain (Meixell and Luoma, 2015). Within this research firms’ CSR activities and reporting are understood to signal their commitments to stakeholders regarding social and environmental issues (Friske et al., 2023). When stakeholders’ sustainability-related expectations are unfulfilled, irresponsible supplier behaviour may be projected onto buying firms, leading to adverse publicity, reputational loss, and costly legal obligations (Bregman et al., 2015). Thus, non-compliance with stakeholders’ requests for sustainability poses a risk to buying firms (Hajmohammad and Vachon, 2016), which this study refers to as supply chain sustainability risk (SCSR). Accordingly, SCSR is defined as “a condition or a potentially occurring event” residing “within a focal firm’s supply chain” which can “provoke harmful stakeholder reactions” (Hofmann et al., 2014, p. 168). During times of global sourcing and ubiquitous information availability, SCSR poses a major challenge to buying firms (Busse et al., 2017).
Previous research suggests that, in general, a success factor for sustainable supply chain management (SSCM) is that buying firms should “reconceptualise who is in the supply chain. Rather than viewing NGOs and the like as adversaries, sustainable supply chains leverage the skills and abilities of these nontraditional chain members” (Pagell and Wu, 2009, p. 52). Hence, an attentive and cooperative stance toward stakeholders is often advisable for firms (Meixell and Luoma, 2015; Wong et al., 2015). To identify, assess, and manage SCSR, firms must understand the differing perspectives, expectations, and values of stakeholders (Wu et al., 2014). Thus, when faced with the lack of visibility in the upstream supply chain, attention to stakeholders may be the strategic direction that firms should also pursue to identify SCSR, seeking to incorporate stakeholders’ SCSR knowledge. Stakeholders vary in numerous ways, including their interests and roles (Wu et al., 2014). Different groups of stakeholders can be interested in the economic, environmental, and social dimensions to different degrees (Meixell and Luoma, 2015). Some stakeholders hope for the firm’s success (e.g., employees and customers), while others may not mind failure (e.g., competitors and media; Hofmann et al., 2014). Therefore, firms should refrain from treating their stakeholders as homogenous aggregates; rather, they should differentiate between them and dedicate specific attention to stakeholder groups (Gualandris et al., 2015).
Regarding stakeholder-OSCM research in the instrumental tradition, considerable attention has been given in OCSM research to the circumstances in which trust, collaboration, mutual investment etc. exist in firm-stakeholder relations, and how these support greater innovation, operational, and financial performance. Empirical studies have often linked supply chain collaboration to performance improvement (Co and Barro, 2009). Managing the relationships among trading partners along the supply chain is an important question that must be addressed in supply chain collaboration (Johnston et al., 2004). A long-term perspective between the buyer and supplier increases the intensity of buyer – supplier coordination (De Toni, 1999), and supplier integration through strategic partnerships “will have a lasting effect on the competitiveness of the entire supply chain” (Choi and Hartley, 1996). Carr and Pearson (1999) also found that strategically managed long-term relationships with key suppliers have a positive impact on the firm’s financial performance.
Supply chain visibility entails trust and collaboration between trading partners. Higher levels of inter-organisational cooperative behaviours, such as shared planning and flexibility in coordinating activities, were found to be strongly linked to the supplier’s trust in the buyer firm (Johnston et al., 2004). Issues of trust and risk can be significantly more important in supply chain relationships because they often involve a higher degree of interdependency between competitors (La Londe, 2002). The premise of trust and collaboration is that trading partners have comparable desires and commitments to collaborate. Morgan and Hunt (1994) argued that commitment and trust produce outcomes that promote efficiency, productivity and effectiveness. Otherwise, it would require an advocate with power and influence over its trading partners to mandate cooperation.
Moving beyond the reasons and stakeholder-OSCM research, considerable research has begun to examine the structural, attitudinal, and informational pre-cursors of successful stakeholder management. For example, research on supply chain visibility emphasises that relationships can be effectively managed only when information is visible across a supply chain and between supply chain partners. Supply chain visibility can be broadly defined as “traceability and transparency of supply chain process” (Tse and Tan, 2012, p. 51). Buying firms often have low supply chain visibility because they possess little knowledge about indirect suppliers or cannot independently verify information about their components or practices (Pagell and Wu, 2009). In particular, fast-moving industries such as the retail and fashion sectors often lack supply chain visibility beyond second-tier suppliers (Roth et al., 2008). Carter et al. (2015) argue that actors in supply chains often lack sufficient knowledge and visibility of their supply chain beyond first-tier suppliers (upstream) and direct customers (downstream), meaning that “what lies beyond the realm of [the] visible range simply emerges” (Carter et al., 2015, p. 90). This visibility boundary poses a severe management problem, since “beyond the visible range, the agent has no choice but to accept what happens there” (Carter et al., 2015, p. 90). Prior research has acknowledged that missing visibility in supply chains is a critical factor in effective supply chain risk management (Durach et al., 2015). Lack of visibility may cause knowledge deficits, loss of control, and distrust, thereby enhancing sustainability and corporate social risks (Busse et al., 2017). The practical challenges of engaging inclusively with a wide range of geographically dispersed stakeholders have been emphasised in research (Siems and Seuring, 2021), with research suggesting that collective stakeholder orientations can help overcome barriers to effective stakeholder involvement (Soundararajan et al., 2019).
@@
Research suggests that firms can adopt various stances toward their stakeholders, ranging from adversarial to welcoming (Pagell and Wu, 2009). Pagell and Wu (2009) observed that leading firms reconceptualise who is in the supply chain, such that they regard not only their direct buyers and suppliers as part of the supply chain, but also other stakeholders. Essentially, they suggest the opening up of firms to these stakeholders. Further studies have also found that some firms leverage the expertise and skills of stakeholders, resulting in better-informed managerial decision-making (Wong et al., 2015). Wong et al. (2015, p. 56) argued that “feedback from [$\dots$] stakeholders represents key resources because (they) sometimes know more about the environmental problems facing part of the supply chains than the focal firm”. Stakeholders can provide assistance, develop policies, engage in evaluation and monitoring, and identify improvement potential in a firm’s upstream supply chain with regard to sustainability (Gualandris et al., 2015; Wong et al., 2015). In the context of SCSR, leading firms have begun to proactively search for valuable information to help them identify their SCSR (or other objectives) by constantly scanning the external environment or conducting regular stakeholder consultations and round tables (Foerstl et al., 2010).
Recent macro-environmental uncertainty and turbulence has encourage OSCM scholarship to explore the role of stakeholders in supporting organisational success in light of these conditions. Indeed, Freeman (1984) himself drew attention to “the quantity and kinds of change that occur in the business environment” (Freeman, 1984, , chapter 5), necessitating a new paradigm of strategic management.
@@
The rapid growth and diversity of stakeholder-related research, both within and beyond OSCM, provide scholars with numerous valuable sources for further reading and inspiration. We have drawn the reader’s attention to some important review articles. For example, Aaltonen and Kujala (2016) review literature on stakeholder management in the project management field; Bhattacharya and Fayezi (2021) review research on stakeholder theory in supply chain management; de Gooyert et al. (2017) review stakeholder management research in the operations research field; Govindan and Bouzon (2018) examine stakeholder-related research in reverse logistics. Sustainable supply chain management has been reviewed extensively by Carter and Liane Easton (2011) and Sarkis et al. (2011). Additionally, there are a number of excellent reviews on the development of stakeholder management outside OSCM, including Laplume et al. (2008). Technology, especially AI and big data analytics, and government policy play a significant role in shaping how firm-stakeholder relationships evolve and are managed, and rapid changes in these in recent years will continue to show that stakeholder relationships play an increasingly salient role in OSCM.
@@
The bullwhip effect refers to a supply chain phenomenon where the variance of customer demand is amplified in the production (or supplier) orders. Figure 7, panel (a) highlights the weekly demand and production for a grocery product. The production order variance is 7.68 times that of demand. This amplification is not uncommon in real supply chains. Panel (b) of figure 7 illustrates the weekly demand and raw material order of a consumable IT product. The variance of the shipments to the assembly factory is 3.58 times that of the demand at the factory. Highly variable demand placed on the upstream echelons of a supply chain creates excess costs. For example, highly variable demand requires extra capacity to meet peak demands; extra inventory is needed to protect service levels; reduced transport efficiency; hiring, firing, and onboarding costs increase.
Figure 7: Examples of real-world bullwhip. Panel a) A UK grocery product, bullwhip ratio = 7.68. Panel b) A US IT consumable product, bullwhip ratio = 3.58.
@@
Control theory has long been one of the most powerful ways of solving linear bullwhip problems. Simon (1952) applied the Laplace transform to a production control problem described by differential equations. Towill (1982) studied an inventory replenishment policy denoted IOBPCS (inventory and order-based production control system) in continuous time using the Laplace transform. John et al. (1994) added a work-in-progress feedback loop to the IOBPCS to create the APIOBPCS model (automatic pipeline, inventory, and order-based production control system). Continuous-time bullwhip measures are commonly determined by the system’s noise bandwidth.
Most supply chains operate on discrete time, where one ordering decision per planning period is made. Vassian (1955) studied a discrete-time replenishment system with z-transforms, finding the inventory optimal order-up-to (OUT) policy. Disney (2001) used the z-transform to study the dynamics of a vendor-managed inventory supply chain. The Fourier transform can be used to understand how supply chains create bullwhip under arbitrary (stationary and non-stationary) demands and (constant) lead times by investigating the system’s frequency response. Dejonckheere et al. (2003) revealed the OUT policy with exponential smoothing and moving average forecasts always creates bullwhip, for all demand patterns and lead times. Li et al. (2014) extended these results, showing the OUT policy with Holt’s forecasts always creates bullwhip, but the OUT policy with damped trend forecasts does not always create bullwhip for certain demand processes. Linear systems can also be studied directly in the time domain and using difference equations and expectation (Hosoda and Disney, 2012). Larger, more complex linear supply chain models can also be analyzed using state space techniques and matrix methods, (Gaalman and Disney, 2006; Wang and Disney, 2017).
@@
Surveys and secondary data can also be used to investigate bullwhip. At an individual product level, bullwhip appears almost inevitable. However, when measured at the company level, perhaps aggregated over a financial reporting period of many months, bullwhip may not exist (Cachon et al., 2007). Bullwhip has also been studied empirically: Fransoo and Wouters (2000) consider how to measure bullwhip in real supply chains (which may have convergent and/or divergent material flows). Disney et al. (2013) report on a project removing bullwhip from a global printer toner supply chain. Meng et al. (2018) improve the forecasting and production planning activities in a UK dairy product manufacturer to reduce bullwhip.
@@
Most bullwhip studies consider high-volume demand where the items are ordered and replenished every planning period. Planning periods range from several hours to a day in grocery supply chains, or weekly (most likely) to monthly in manufacturing environments. Hedenstierna and Disney (2018) investigate the consequences of different planning cycle lengths; longer cycles have a natural pooling effect that reduces demand variance. With no ordering (capacity-related) costs, the OUT policy is the optimal replenishment policy for minimising inventory holding and backlog costs (Karlin and Scarf, 1958). When the objective function includes bullwhip-related (capacity-related) costs, the OUT policy is no longer optimal, Boute et al. (2022); the dual base stock (DBS) policy is known to be the optimal policy (Boute and Van Mieghem, 2015; Gijsbrechts et al., 2024). However, the DBS policy is not a demand replacement policy. The DBS policy has a region of inaction where demand is not fully replaced and only $k$ units (where $k$ is the nominal capacity without overtime) are ordered. Generally, only numerical solutions for the two base stock levels in the DBS policy can be obtained (Gijsbrechts et al., 2024).
The high-performing linear proportional (POUT) policy, grounded in control theory, is amenable to analytical study. The POUT policy generalizes the OUT policy by incorporating two proportional feedback controllers into the policy, an established technique in the design of hardware systems. The feedback controllers in the POUT policy regulate the speed at which discrepancies in inventory and WIP levels are corrected (Disney and Towill, 2003). While the POUT policy requires careful set-up to avoid instability (Disney, 2008), it is the optimal linear quadratic regulator for minimising costs proportional to the variance of order and inventory (Wang and Disney, 2016). When piece-wise linear convex inventory costs and capacity costs associated with an installed capacity and overtime are present Boute et al. (2022) show the POUT performs well compared to the optimal non-linear DBS policy.
@@
Most bullwhip studies assume demand is random. The simplest random demand is the independent and identically distributed (i.i.d.) demand. Forecasted with minimum mean squared error (MMSE) forecasts, i.i.d. demand produces a bullwhip ratio of unity, as the OUT policy acts as a pass-on-orders policy.
@@
Closed loop supply chains are popular areas for bullwhip studies (Ponte et al., 2022). Marketplace returns can have a positive effect on the production of new items. However, care must be taken as it is not always economical to remanufacture everything returned, even when the remanufacturing cost is less than the cost to produce new items (Hosoda et al., 2021).
Issues with global dual sourcing have also been studied from a bullwhip perspective. The tailored-base-surge allocation scheme (Allon and Van Mieghem, 2010), places large constant orders to the offshore supplier, whereas the nearshore supplier satisfies peak demands with a shorter lead time. Boute et al. (2022) showed that local SpeedFactories could maintain tighter inventory control than a single offshore supplier, facilitating the relocation of production back home before reaching price parity.
Emerging bullwhip research topics
There are a number of critical areas related to the bullwhip effect that are relatively underexplored. For example, understanding the role of artificial intelligence/machine learning (AI/ML) in forecasting demand and regulating production and distribution. There is a lot of industrial interest in using AI/ML to forecast demand. Presumably, the large and diverse datasets that can be used as input data will lead to more accurate forecasts. However, in cases where we know the optimal replenishment policy for a given cost function (for example, the OUT policy is the optimal linear inventory cost policy for a given set of forecasts), ML/AI may not be able to improve upon current best practice.
Geopolitical issues such as tariffs, trade barriers, conflicts, and sanctions are altering supply chains and trading relationships. As companies restructure their supply chains, there is an opportunity to design out bullwhip problems by shortening lead times, removing supply chain echelons, re-shoring, and dual sourcing. The impact of bullwhip on the carbon emissions and other climate consequences generated by supply chains is also an important area for future bullwhip research.